# Paloma[Uncaptioned image] : A Benchmark for Evaluating Language Model Fit

### v1

@@ 4.2 Scaling Improves Domain Fit Unequally
We return to our initial question: Does rising performance lift all domains? That is, does the sign of scaling trends observed in previous work (Kaplan et al., 2020; Hoffmann et al., 2022) hold across all domains? And if so, do some domains still capture most of the improvement while others stagnate?
4.2.1 Scaling Tokens Seen
@@
In Figure 5, We study the impact of increased training on domain fit. We make use of the finding that the logarithms of loss and tokens seen trend linearly Kaplan et al. (2020), and make an estimate of improvement based on the slope between two empirical observations, with some inital and final number of tokens seen by checkpoints of a model $\theta$:
[…]
@@
4.2.2 Scaling Model Parameters
While the 6 baseline models that we pretrain ourselves are all 1B parameter models, we can use models of varying sizes from the Pythia model suite (Biderman et al., 2023) to examine the impact of scaling model parameters on domain fit. As we note in §2.3, these models are not controlled for contamination but they do address all of our other guidelines.
@@ 4.2.2 Scaling Model Parameters
[…]
@@
4.3 Common Vocabulary Types Dominate Perplexity, Others Have Inverse Scaling
(a) Mean loss per type
(b) Cumulative proportion of total loss per type
Figure 9: Mean and total loss per vocabulary type, i.e., specific strings represented in the tokenizer vocabulary. While high-frequency types (which have low IDs) tend to have a low average loss as shown by a log-linear regression (a), they contribute a substantial part of the total loss, simply by virtue of their frequent occurrence in the data (b). The figure shows the distributions for Pythia-7B on C4-100-domains, but the overall picture is consistent for different models and sources.
@@ 4.3 Common Vocabulary Types Dominate Perplexity, Others Have Inverse Scaling
So far we have examined perplexity aggregated over tokens. As introduced in §3, another approach is to measure surprise over occurrences of specific strings. In Paloma we measure average likelihood per vocabulary type, i.e., the strings that are represented in the vocabulary of a model, in contrast to occurrences of these strings in some corpus, called tokens.
How much do types contribute to the likelihoods aggregated per token in perplexity? To answer this question, we start by analyzing the total loss mass added by types, as a function of their IDs. Given how the GPTNeoX-20B tokenizer was trained (Sennrich et al., 2016; Black et al., 2022), smaller IDs correspond to more frequent types in the tokenizer training data, and we find an overall moderate to strong correlation between IDs and frequencies in the evaluation data of Paloma as well (Pearson’s $r$ averaged across domains: –0.522$\pm$0.087). Crucially, frequency has a strong impact on the total loss mass associated with individual types: while the average loss is lower for the high-frequency types (Figure 8(a)), the total loss is higher, resulting in a situation where 5% of the types already cover roughly 50% of the overall perplexity (Figure 8(b)). Thus, perplexity is strongly influenced by a relatively small set of high-frequency types.
@@
What does it mean specifically if Pythia-1B has a lower average loss on a specific type than Pythia-7B? Figure 12 shows, for each of the 18 sources, the training dynamics of an example type for which Pythia-1B is better than Pythia-7B after convergence. As can be seen, there are various patterns: sometimes there is a constant gap between the two models, with Pythia-1B being better from the very beginning (e.g., Boat in Falcon Refinedweb); sometimes Pythia-1B has a constant loss while Pythia-7B is getting worse over time (e.g., schedule in Dolma); sometimes Pythia-7B has a constant loss while Pythia-1B is getting better over time (e.g., exchanged in The Pile); finally, sometimes Pythia-1B is decreasing its loss while Pythia-7B is increasing its loss over time (e.g., BR in C4). Especially the last pattern bears a resemblance with inverse scaling effects that characterize other aspects of LM behavior, where the performance gets worse rather than better with larger models (Mckenzie et al., 2023). We are not aware of prior work describing the kind of type-level inverse scaling that we observe in this analysis.
@@
5 Related Work
Previous fine-grained perplexity corpora
Paloma is inspired by and incorporates previous work that curates corpora with marked domains. The Pile (Gao et al., 2020) is an early public pretraining corpus that provides provenance-based subdivisions of data rather than homogeneous webscraped data. Their 22 domains are included among the 585 domains in Paloma. More recent works curate datasets with on the order of hundreds of domains (Reid et al., 2022; Chronopoulou et al., 2022; Li et al., 2022), which we include in part or full in Paloma. Likewise we include efforts that curate text from specific dialects (Blodgett et al., 2016; Greenbaum & Nelson, 1996).
Previous LLM benchmarks
Language modeling is a field of research with a long history and many classic benchmarks (Baker et al., 1983, Chelba et al., 2013, Merity et al., 2016, and Marcus et al., 1999, inter alia). With the more recent success of language modeling as a pretraining step, benchmarks have shifted towards downstream task evaluations rather than perplexity metrics. Perhaps the most notable exception is The Pile (Gao et al., 2020) which, in addition to being a pretraining corpus, is explicitly framed as a benchmark and provides a detailed specification of the formatting of inputs for inference to compute perplexity. Where recent LMs report perplexity beyond the model’s own training distribution (Rae et al., 2021; Hoffmann et al., 2022), it is almost always on The Pile. We adopt their inference format and add additional controls for contamination, data order, and tokenization to our benchmark (see Table 1). Other efforts focus on comprehensive LM evaluation predominantly using downstream tasks (Liang et al., 2022; Gao et al., 2021). We employ one such open source evaluation tool, Catwalk (Groeneveld et al., 2023), [^6] https://github.com/allenai/catwalk to compute our metrics. Moreover, we see our work as complementary to efforts in downstream evaluation, as previous work disagrees whether perplexity evaluations are predictive of downstream performance (Liu et al., 2022; Tay et al., 2021; Ganguli et al., 2022; Xia et al., 2022). In fact, we hope that our benchmark can provide well controlled perplexity results for further study of this question.
6 Conclusion
We believe that evaluations of LM fit provide an important view of language model performance that has been neglected in recent LM research and development. Perplexity cannot be naïvely applied to language modeling at this scale due to challenges such as benchmark contamination. However, these obstacles are worth overcoming as perplexity offers several advantages not afforded by downstream evaluations. Instead of constructing tasks from scratch, we can rely on the ecological validity of real-world data drawn from known sources. By figuring out the best ways to evaluate fit of a model to a collection of documents, we create an interface for other fields to easily compose evaluations for language models. Researchers in other fields need not understand the architectures of such models to collect a corpus of text representing domains of interest that LM researchers would not know to consider. Once a significant data source is identified, such as an online community, evaluations can be updated over time by simply scraping more data, unlike downstream tasks where expensive annotation would be required.
Paloma advances the possibilities for evaluation of LM fit by providing finely categorized domains of text and controls to mitigate confounders of our metrics. We hope that, along with the baseline models that we train for this work, submissions will begin to fill in the space of possible LM configurations over choices such as data composition, data size, parameter count, and learning rate schedule. By encouraging standardization within this space we increase the density of comparisons that can be made between models. With greater resolution of evaluations, we can increase the resolution of our understanding of language model training dynamics.
7 Limitations
The largest limitation of Paloma is that we elect to focus just on the language modeling of English and code data. We select this scope as most current LMs also focus on theses types of data. However, we strongly encourage future work to explore how language model fit to fine-grained domains behaves within and across other languages. We also wish to note that while we measure performance on more fine-grained domains, our choice of metrics may not reflect what is valued by all the communities who produce language in these domains (Diaz & Madaio, 2023). Nevertheless, we think that examining discrepancies in existing metrics over domains will lead to a deeper understanding of language modeling dynamics, which can then illuminate the gaps in existing approaches to evaluation. In this we follow Holtzman et al. (2023) and McCoy et al. (2023) by aiming to examine model behaviors, regardless of whether those behaviors are desirable or not to humans.
We also caution against hasty interpretation of LM fit to any particular domain as being indicative of alignment of a model to features that a human might consider salient about these domains. For instance we find that when just examining perplexity, results on the 3 fringe datasets are tightly related to average document lengths, with the short tweet-like posts in Gab Corpus receiving high perplexities while the long connected threads of posts in 4chan Corpus and Manosphere Corpus provide greater context and lower perplexity. At this level of aggregation, differences in surprise between these domains likely have little to do with model fit to specific types of toxicity. In our case study in §4.3, we demonstrate that often it is more appropriate to decompose measures of surprise over specific strings within a corpus, rather than aggregating over all text in a domain. We hope that by surfacing the average likelihoods of specific strings in the vocabulary, Paloma can enable future work on metrics that better measure the fit of models to the features of language in specific domains that humans find most salient.
Another set of limitations arises from our use of documents as a fundamental unit of data. When subsampling although we balance the number of tokens used to represent each domain, we still sample documents until that target token count is reached. Concretely this means that some domains, especially books, are represented by only dozens of documents, which likely does not capture the full distribution of the domain as well many smaller documents might. This also impacts our decontamination approach, since we remove whole documents that have any paragraph marked as contaminated to avoid mangling documents by excising individual paragraphs. Such an approach tends to disproportionately remove long documents that are frequently quoted, which may include seminal works along the lines of Martin Luther King’s “I Have a Dream” speech that actually deployed models should be familiar with. The purpose of Paloma, however, is to enable controlled research on the science of language modeling, but production models should likely use caution in applying this decontamination technique.
As we note in §2.4.1, our measures of efficiency aim to consider cost agnostic of hardware. It is important to acknowledge that it is not possible to completely separate hardware from abstract notions of cost. For instance, models that utilize greater parallelism through sparse combinations of predictions from experts will appear to have parameter counts equal to the sum of experts even though those parameters can be trained in far less wall-clock time on multiple devices. Further though we record when submissions have run multiple epochs over data, this does not fully disentangle data and compute efficiency. That is datasets such as The Pile can include intentional oversampling of data within a single epoch, and other datasets include multiple different prepossessed versions of the same underlying data (for instance including both C4 and some other version on Common Crawl data).
Finally we would like to acknowledge that we are unable to rehost The Pile and ICE data in Paloma for easy access due to restrictions on these datasets. Permission to access ICE can be arranged through contacting the original authors. [^7] https://www.ice-corpora.uzh.ch/en/access.html
Acknowledgements
Appendix A Evaluation Data Source Details
@@
Initially the pretraining corpus used by Raffel et al. (2019) and later released in Dodge et al. (2021), C4 has become one of the most commonly used pretraining corpora and is often included in more recently curated corpora. It uses a single April 2019 Common Crawl scrape to source webtext. This is filtered to remove text that is not classified as English as well as heuristics to remove text that is not natural language and a blocklist of profane keywords. We sample from the validation split of this "cleaned" corpus to measure model fit to webtext from a single temporal slice of scraping with baseline preprocessing. This source has no marked domains.
@@
Chung et al. (2023) release a dataset with same the methods used in C4 but scale up to all Common Crawl scrapes up to August 2022 and include 107 classified languages. As the scope of the present work is the evaluation of English language models we sample only from the validation split of the English portion of the data. This allow us to measure the fit of models to scraped webtext with heterogeneous temporality. This source has no marked domains.
The Pile
Gao et al. (2020) curate a pretraining corpus from 22 domains in one of the first large open corpora to include mostly non-webscraped text, such as archives of novels or academic papers. It is also explicitly framed as a language modeling benchmark with instructions for standardized evaluations on the validation and test sets, and several open source models have been trained on it (Wang & Komatsuzaki, 2021; Black et al., 2022; Biderman et al., 2023). We sample its 22 domains (see Table 5) to evaluate fit to coarse-grained, mostly non-webscraped domains.
Table 5: Domains in The Pile
@@
We include these two benchmarks as they have seen the most consistent evaluation on large LMs. Wikitext-103 (Merity et al., 2016) consists Wikipedia articles marked “Good” and “Featured” and was used in the evaluation of GPT-2 (Radford et al., 2019), Gopher (Rae et al., 2021), and Chinchilla (Hoffmann et al., 2022). Penn Treebank (Marcus et al., 1999) consists of 1989 Wall Street Journal articles originally annotated for linguistic structure. GPT-2 (Radford et al., 2019) and GPT-3 (Brown et al., 2020) omit these annotations and evaluate perplexity on the underlying text. We sample the same version of the benchmark, which is hosted by Nunes (2020). As was standard practice at the time the benchmark is pretokenized and uncommon words are replaced with a special unknown token; we opt not to detokenize this data as we find contemporary LMs are often able to achieve comparable performance to the GPT-3 SOTA without this. These two sources have no marked domains.
@@
Together Computer (2023) reproduce a pretraining corpus following the data mixture of LLaMA (Touvron et al., 2023), which combines curated sources and webscraped text similarly to The Pile but with a much greater portion of scraped data as has become customary in recent pretraining corpora. This dataset is used to train RedPajama-INCITE (Together Computer, 2023), one of the few models with both checkpoints and data publicly available. We sample their 7 domains (see Table 6).
@@
Included in the training of the Falcon models (Almazrouei et al., 2023), Penedo et al. (2023) collect a corpus of English sampled from all Common Crawl scrapes until June 2023. While we include other Common Crawl based corpora, this one has a higher duplication removal rate than previous corpora. They also claim to have more neutral filters that rely on simple interpretable heuristics and only blocklist adult content by URLs. We sample this to examine how differences in filtering scraped data influence perplexity evaluations. This source has no marked domains.
@@
Soldaini et al. (2023) curate a corpus from Common Crawl, Wikipedia, books, academic papers, code repositories, and Reddit—domains similar to those used to train most contemporary LLMs. They release the code used to collect and process this data which in combination with the corpus serve as a set of scientific artifacts to support broader participation in research on pretraining data. We sample from held out splits of each of these domains (see Table 7) to provide corresponding evaluations for these artifacts.
@@
Reid et al. (2022) collect academic papers from S2ORC (Lo et al., 2020) and organize them into a two level hierarchy by academic field categories. Top-level domains, such as Computer Science, are already provided in S2ORC using top-level disciplines from the Microsoft Academic Graph (Shen et al., 2018), while subdomains are identified by a paper’s arXiv category, such as the subdomain Computation and Language within Computer Science. As academic papers are a common source for pretraining and a domain for downstream use, we sample from this corpus to measure fine-grained fit to different academic disciplines. We sample both their top-level domains and lower-level subdomains, as our definition of domain accepts that domains may overlap. Also note that while the M2D2 paper only reports 106 domains and subdomains of S2ORC data, we find that there are actually 167 domains and subdomains (see Table 8) marked in their final corpus. Unfortunately the original collection concatenates together all papers, making it impossible to recover document boundaries. We resort instead to sampling a given number of tokens from the beginning of the concatenated sequences as one long pseudo-document, relying on the random shuffling of the original data before concatenation.
@@
Reid et al. (2022) also collect Wikipedia articles and organize them by the top two levels of hierarchy from the Wikipedia ontology. We sample from this source, as the Wikipedia ontology provides some of the largest scale human categorization of domains of text available on a data source almost always included in pretraining corpora. This time we find that their corpus contains just 49 marked domains or subdomains (see Table 9), rather than the 60 mentioned in the paper. Again the original collection concatenates articles together, so we sample a given number of tokens from the beginning of this concatenated sequence.
@@
Chronopoulou et al. (2022) collect C4-100-domains comprising all the text from 100 internet domains with the most pages in C4. We sample from each of the 100 domains (see Table 10) to explore the relationship between how well represented and how surprising a domain is. The original collection removes documents smaller than 200 whitespace separated tokens, leading the domain with the 3rd most pages (do5.b00kmedia.ru) to be completely empty. Only three other domains have less data than the 100 thousand tokens per split that we aim for.
@@
Using the Reddit data collected in Dolma (Soldaini et al., 2023), we organize a new corpus of the top 100 subreddits (community forums within the messageboard) ranked by number of posts in the Dolma data (see Table 11). In Dolma Reddit posts are each separate documents, without any linearization of conversational threads. Though this prevents the assessment of model fit to dialogue, it still allows evaluation across these many domains of social media text. The Dolma Reddit data also filters out comments shorter than 500 characters and submissions (i.e., original posts) shorter than 400 characters. We sample these subreddits to capture domains as they are self-organized and self-identified by online communities.
@@
Using code repository data from The Stack (Kocetkov et al., 2022) as it is contained in Dolma (Soldaini et al., 2023), we collect a new corpus of balanced samples of the top one hundred programming languages by number of tokens (see Table 12). While code data differs greatly from natural language, complicating the interpretation of perplexity analysis, we nevertheless wish to add evaluations to cover this common data source for LLMs.
@@
ICE
Local research teams following guidelines established in Greenbaum & Nelson (1996) collected corpora of English from Canada, East Africa (Kenya & Tanzania), Hong Kong, India, Ireland, Jamaica, Philippines, Singapore, and the USA. Each of these samples of English from around the world is further split into a written and transcribed spoken corpus, except for USA which only has written data (see Table 13). We follow HELM (Liang et al., 2022) in utilizing this corpus to measure disparate performance between these dialects. To permit comparability to HELM, we follow the same preprocessing which leaves in some XML-style tags marking phenomena such as speaker turns.
Table 13: Domains in ICE
@@
Blodgett et al. (2016) create a pair of corpora representing African-American and White-aligned English using a statistical model with distant supervision from geolocation and demographic census statistics. We follow the reproduction of this dataset used in HELM (Liang et al., 2022), but we fix an error in loading escaped sequences of the data that, among other issues, renders emojis as literal hexadecimal bytes. Our reproduction is not able to sample the same documents, but is otherwise identical. We sample these corpora to examine disparities in performance on minoritized dialects (see Table 14).
@@
Ribeiro et al. (2021) curate a corpus of texts spanning 2006 to 2019 scrapped from 9 forums sharing a masculinist ideology: 8 independent message boards as well as 56 subreddits on Reddit. Using a toxicity classifier and lexicon-based misogyny metric, they find an increase in toxicity and hate over time to levels far above mainstream Reddit and comparable to 4chan Corpus. We sample this corpus to measure fit to a discourse with a specific variety of toxicity focused on hate towards women. Moreover we intend this to exemplify how domain expertise allows the manual curation of a corpus to represent a whole discourse using known relationships between sources. The original data already linearizes the posts into a sequential thread, which we concatenate together with post authors prepended to posts. Though this datasets marks 9 domains (see Table 15), we opt to treat this whole source as a single domain for the present analysis and thus do not perform a stratified sample of these domains.
@@
Zannettou et al. (2018) scrape posts from August 2016 and January 2018 on Gab, an alt-right focused Twitter alternative founded in 2016. The platform emphasizes freedom of speech and minimal moderation, with notable users joining after being banned from mainstream social media. The authors find that Gab Corpus measures higher than Twitter but lower than 4chan Corpus on a lexicon of hate words. We sample this corpus to measure fit to low moderation social media. We treat posts as independent documents, rather than attempting to reconstruct connected subgraphs of posts replying to other posts. This source has no marked domains.

### v2

@@ D.1 Scaling Improves Domain Fit Unequally
We return to the question, does rising performance lift all domains? That is, does the sign of scaling trends observed in previous work [Kaplan et al., 2020, Hoffmann et al., 2022] hold across all domains? And if so, do some domains still capture most of the improvement while others stagnate?
D.1.1 Scaling Tokens Seen
@@
In Figure 6, we study the impact of increased training on domain fit. We make use of the finding that the logarithms of loss and tokens seen trend linearly Kaplan et al. [2020], and make an estimate of improvement based on the slope between two empirical observations of perplexity (ppl), with some initial and final number of tokens, $i$ and $f$, seen by checkpoints of a model $\theta$:
$\Delta_{t}(i,f)=\frac{\text{ln}(\text{ln}(\text{ppl}(\theta_{i})))-\text{ln}(\text{ln}(\text{ppl}(\theta_{f})))}{\text{log}_{10}(f)-\text{log}_{10}(i)}$
@@
D.1.2 Scaling Model Parameters
While the 6 baseline models that we pretrain ourselves are all 1B parameter models, we can use models of varying sizes from the Pythia model suite [Biderman et al., 2023] to examine the impact of scaling model parameters on domain fit. As we note in §G, these models are not controlled for contamination but they do address all of our other guidelines.
@@ D.1.2 Scaling Model Parameters
In Figure 9 we perform the same analysis of improvement in log loss as before but this time with respect to log increase in non-embedding parameters, $\Delta_{p}(i,f)$. Specifically we plot $\Delta_{p}(85M,805M)$ and $\Delta_{p}(805M,6.4B)$ for the non-embedding parameter counts corresponding to the 160M, 1B, and 7B model sizes for each domain in ascending order per pair of models compared. This time scaling does universally result in improvements. However, the rate of improvement varies greatly from domain to domain. Examining the median difference in improvement between most and least improved domains shows 2.02$\times$ improvement for the most improved domain, a similar gap to that seen on increases in tokens seen. Again, we stress that unequal improvement is not necessarily problematic, but rather it helps identify outlier domains that follow different scaling trends than the majority of the data. We offer examples of most and least improved domains with respect to increase in model size in the Appendix in Figure 14.
@@
D.2 Common Vocabulary Types Dominate Perplexity, Others Have Inverse Scaling
@@
Previously in §4.2 we noted that few vocabulary types account for most of the loss measured in perplexity. Now we continue to explore the dynamics of average likelihood per vocabulary type, i.e., the strings that are represented in the vocabulary of a model (Appendix B).
@@
What does it mean specifically if Pythia-1B has a lower average loss on a specific type than Pythia-7B? Figure 12 shows, for each of the 16 sources, the training dynamics of an example type for which Pythia-1B is better than Pythia-7B after convergence. As can be seen, there are various patterns: sometimes there is a constant gap between the two models, with Pythia-1B being better from the very beginning (e.g., Boat in Falcon Refinedweb); sometimes Pythia-1B has a constant loss while Pythia-7B is getting worse over time (e.g., schedule in Dolma); sometimes Pythia-7B has a constant loss while Pythia-1B is getting better over time (e.g., exchanged in The Pile); finally, sometimes Pythia-1B is decreasing its loss while Pythia-7B is increasing its loss over time (e.g., BR in C4). Especially the last pattern bears a resemblance with inverse scaling effects that characterize other aspects of LM behavior, where the performance gets worse rather than better with larger models [Mckenzie et al., 2023]. We are not aware of prior work describing the kind of type-level inverse scaling that we observe in this analysis.
@@
Appendix E Evaluation Data Source Details
Purpose Source Reference Description
Standard language modeling benchmarks C4 Raffel et al. [2019] via Dodge et al. [2021] Standard contemporary LM pretraining corpus automatically filtered from the April 2019 Common Crawl scrape
mC4-en Chung et al. [2023] The English language portion of a pretraining corpus automatically filtered from 71 Common Crawl scrapes
WikiText-103 Merity et al. [2016] A standard collection of verified “Good” and “Featured” articles on Wikipedia
Penn Treebank Marcus et al. [1999] via Nunes [2020] Classic Wall Street Journal benchmark with linguistic structure annotations omitted
RedPajama Together Computer [2023] A publicly available reproduction of the LLaMA [Touvron et al., 2023] pretraining source mixture, combining large amounts of webscraped text with smaller curated sources
Falcon Refinedweb Penedo et al. [2023] A corpus of English sampled from all Common Crawl scrapes until June 2023, more aggressively filtered and deduplicated than C4 and mC4-en
Dolma Soldaini et al. [2024] A three trillion token corpus that samples sources commonly used to train LMs in order to enable open research on pretraining data
Fine-grained domain benchmarks M2D2 S2ORC Reid et al. [2022] Papers from Semantic Scholar grouped by hierarchical academic field categories
M2D2 Wikipedia Reid et al. [2022] Wikipedia articles grouped by hierarchical categories in the Wikipedia ontology
C4-100-domains Chronopoulou et al. [2022] Balanced samples of the top 100 URL domains in C4 as measured by page count
Dolma-100-subreddits Soldaini et al. [2024] Balanced samples of the top 100 subreddits by number of posts, sourced from the Dolma Reddit subset
Dolma-100-programming-languages Kocetkov et al. [2022] via Soldaini et al. [2024] Balanced samples of the top 100 programming languages by number of tokens, sourced from the Dolma Stack subset
Communities disparities TwitterAAE Blodgett et al. [2016] via Liang et al. [2022] Balanced sets of tweets classified as African American or White aligned English
Fringe sources previously studied for problematic discourse Manosphere Corpus Ribeiro et al. [2021] 9 forums where a set of related masculinist ideologies developed over the 2000s and 2010s
Gab Corpus Zannettou et al. [2018] Data from 2016-2018 from an alt-right, free-speech-oriented social media platform shown to contain more hate speech than mainstream platforms
4chan Corpus Papasavva et al. [2020] Data from 2016-2019 from a politics subforum of an anonymity-focused forum found to contain among the highest rates of toxic content
Table 5: Descriptions of the 16 data sources sampled to create language modeling evaluations in Paloma. These are grouped by their purposes for inclusion (§2).
In Table 5 we summarize each data source. All sources are existing research datasets and thus we believe our of these datasets for an evaluation benchmark is consistent with their intended use. These sources are permissively licensed and thus we are able to rehost them. Also note that we make no attempt to remove personally identifiable information (PII) beyond any filtering applied by these original datasets. As we rehost only small subsamples of these datasets and the full datasets are also publicly available, any malicious use of these datasets would simply bypass any additional filtering we could do by using the original datasets. Also our subsampling is random and thus does not make it easier for malicious use to aggregate PII.
In the rest of this section we provide details of our use of each source and list all domains if any in each source.
@@ Appendix E Evaluation Data Source Details
Initially the pretraining corpus used by Raffel et al. [2019] and later released in Dodge et al. [2021], C4 has become one of the most commonly used pretraining corpora and is often included in more recently curated corpora. It uses a single April 2019 Common Crawl scrape to source webtext. This is filtered to remove text that is not classified as English as well as heuristics to remove text that is not natural language and a blocklist of profane keywords. We sample from the validation split of this "cleaned" corpus to measure model fit to webtext from a single temporal slice of scraping with baseline preprocessing. This source has no marked domains.
@@
Chung et al. [2023] release a dataset with same the methods used in C4 but scale up to all Common Crawl scrapes up to August 2022 and include 107 classified languages. As the scope of the present work is the evaluation of English language models we sample only from the validation split of the English portion of the data. This allow us to measure the fit of models to scraped webtext with heterogeneous temporality. This source has no marked domains.
@@
We include these two benchmarks as they have seen the most consistent evaluation on large LMs. WikiText-103 [Merity et al., 2016] consists Wikipedia articles marked “Good” and “Featured” and was used in the evaluation of GPT-2 [Radford et al., 2019], Gopher [Rae et al., 2021], and Chinchilla [Hoffmann et al., 2022]. Penn Treebank [Marcus et al., 1999] consists of 1989 Wall Street Journal articles originally annotated for linguistic structure. GPT-2 [Radford et al., 2019] and GPT-3 [Brown et al., 2020] omit these annotations and evaluate perplexity on the underlying text. We sample the same version of the benchmark, which is hosted by Nunes [2020]. As was standard practice at the time the benchmark is pretokenized and uncommon words are replaced with a special unknown token; we opt not to detokenize this data as we find contemporary LMs are often able to achieve comparable performance to the GPT-3 SOTA without this. These two sources have no marked domains.
@@
Together Computer [2023] reproduce a pretraining corpus following the data mixture of LLaMA [Touvron et al., 2023], which combines curated sources and webscraped text similarly to The Pile but with a much greater portion of scraped data as has become customary in recent pretraining corpora. This dataset is used to train RedPajama-INCITE [Together Computer, 2023], one of the few models with both checkpoints and data publicly available. We sample their 7 domains (see Table 6).
@@
Included in the training of the Falcon models [Almazrouei et al., 2023], Penedo et al. [2023] collect a corpus of English sampled from all Common Crawl scrapes until June 2023. While we include other Common Crawl based corpora, this one has a higher duplication removal rate than previous corpora. They also claim to have more neutral filters that rely on simple interpretable heuristics and only blocklist adult content by URLs. We sample this to examine how differences in filtering scraped data influence perplexity evaluations. This source has no marked domains.
@@
Soldaini et al. [2024] curate a corpus from Common Crawl, Wikipedia, books, academic papers, code repositories, and Reddit—domains similar to those used to train most contemporary LLMs. They release the code used to collect and process this data which in combination with the corpus serve as a set of scientific artifacts to support broader participation in research on pretraining data. We sample from held out splits of each of these domains (see Table 7) to provide corresponding evaluations for these artifacts.
@@
Reid et al. [2022] collect academic papers from S2ORC [Lo et al., 2020] and organize them into a two level hierarchy by academic field categories. Top-level domains, such as Computer Science, are already provided in S2ORC using top-level disciplines from the Microsoft Academic Graph [Shen et al., 2018], while subdomains are identified by a paper’s arXiv category, such as the subdomain Computation and Language within Computer Science. As academic papers are a common source for pretraining and a domain for downstream use, we sample from this corpus to measure fine-grained fit to different academic disciplines. We sample both their top-level domains and lower-level subdomains, as our definition of domain accepts that domains may overlap. Also note that while the M2D2 paper only reports 106 domains and subdomains of S2ORC data, we find that there are actually 167 domains and subdomains (see Table 8) marked in their final corpus. Unfortunately the original collection concatenates together all papers, making it impossible to recover document boundaries. We resort instead to sampling a given number of tokens from the beginning of the concatenated sequences as one long pseudo-document, relying on the random shuffling of the original data before concatenation.
@@
Reid et al. [2022] also collect Wikipedia articles and organize them by the top two levels of hierarchy from the Wikipedia ontology. We sample from this source, as the Wikipedia ontology provides some of the largest scale human categorization of domains of text available on a data source almost always included in pretraining corpora. This time we find that their corpus contains just 49 marked domains or subdomains (see Table 9), rather than the 60 mentioned in the paper. Again the original collection concatenates articles together, so we sample a given number of tokens from the beginning of this concatenated sequence.
@@
Chronopoulou et al. [2022] collect C4-100-domains comprising all the text from 100 internet domains with the most pages in C4. We sample from each of the 100 domains (see Table 10) to explore the relationship between how well represented and how surprising a domain is. The original collection removes documents smaller than 200 whitespace separated tokens, leading the domain with the 3rd most pages (do5.b00kmedia.ru) to be completely empty. Only three other domains have less data than the 100 thousand tokens per split that we aim for.
@@
Using the Reddit data collected in Dolma [Soldaini et al., 2024], we organize a new corpus of the top 100 subreddits (community forums within the messageboard) ranked by number of posts in the Dolma data (see Table 11). In Dolma Reddit posts are each separate documents, without any linearization of conversational threads. Though this prevents the assessment of model fit to dialogue, it still allows evaluation across these many domains of social media text. The Dolma Reddit data also filters out comments shorter than 500 characters and submissions (i.e., original posts) shorter than 400 characters. We sample these subreddits to capture domains as they are self-organized and self-identified by online communities.
@@
Using code repository data from The Stack [Kocetkov et al., 2022] as it is contained in Dolma [Soldaini et al., 2024], we collect a new corpus of balanced samples of the top one hundred programming languages by number of tokens (see Table 12). Dolma uses an already near-deduplicated version of The Stack, filters data related extensions (e.g., JSON and CSV) and repetitive preambles, and applies quality heuristics (e.g., removing repos with few stars). While code data differs greatly from natural language, complicating the interpretation of perplexity analysis, we nevertheless wish to add evaluations to cover this common data source for LLMs.
@@
Blodgett et al. [2016] create a pair of corpora representing African-American and White-aligned English using a statistical model with distant supervision from geolocation and demographic census statistics. We follow the reproduction of this dataset used in HELM [Liang et al., 2022], but we fix an error in loading escaped sequences of the data that, among other issues, renders emojis as literal hexadecimal bytes. Our reproduction is not able to sample the same documents, but is otherwise identical. We sample these corpora to examine disparities in performance on minoritized dialects (see Table 13).
@@
Ribeiro et al. [2021] curate a corpus of texts spanning 2006 to 2019 scrapped from 9 forums sharing a masculinist ideology: 8 independent message boards as well as 56 subreddits on Reddit. Using a toxicity classifier and lexicon-based misogyny metric, they find an increase in toxicity and hate over time to levels far above mainstream Reddit and comparable to 4chan Corpus. We sample this corpus to measure fit to a discourse with a specific variety of toxicity focused on hate towards women. Moreover we intend this to exemplify how domain expertise allows the manual curation of a corpus to represent a whole discourse using known relationships between sources. The original data already linearizes the posts into a sequential thread, which we concatenate together with post authors prepended to posts. Though this datasets marks 9 domains (see Table 14), we opt to treat this whole source as a single domain for the present analysis and thus do not perform a stratified sample of these domains.
@@
Zannettou et al. [2018] scrape posts from August 2016 and January 2018 on Gab, an alt-right focused Twitter alternative founded in 2016. The platform emphasizes freedom of speech and minimal moderation, with notable users joining after being banned from mainstream social media. The authors find that Gab Corpus measures higher than Twitter but lower than 4chan Corpus on a lexicon of hate words. We sample this corpus to measure fit to low moderation social media. We treat posts as independent documents, rather than attempting to reconstruct connected subgraphs of posts replying to other posts. This source has no marked domains.